OpenStudy (anonymous):

Solve -4(x + 1) - 3 = -3(x - 4).

OpenStudy (anonymous):

-4x-4-3=-3x+12 -4x-7=-3x+12 -x=19 x=-19

OpenStudy (anonymous):

thank you